# Rapid formation of a very massive star $>\!\!50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

### v1

@@ 5.2 Observational context for VMS formation at high redshift
Observing massive, compact, and low-metallicity star clusters around the time of core collapse remains extremely challenging. However, high-redshift galaxies are widely expected to form such clusters efficiently due to their high gas fractions and low metallicities (e.g. Tacconi et al., 2010), as well as their clumpy substructures that favor gravitational fragmentation and rapid star formation (e.g. Devereaux et al., 2023). These conditions closely mirror those adopted in our simulations, reinforcing the plausibility of early VMS formation via runaway stellar collisions.
[…]
@@
From a theoretical standpoint, the Toomre instability criterion implies that the typical cluster size should scale inversely with the square root of the gas surface density, i.e., $R_{h}\propto\Sigma_{\mathrm{gas}}^{-1/2}$ (Toomre, 1964). This suggests that star clusters formed in gas-rich, high-redshift galaxies should be significantly more compact than their local counterparts—a trend also supported by simulations and nearby observations (Choksi & Kruijssen, 2021). Reinforcing this, Adamo et al. (2024) recently reported several compact clusters at $z\sim 10$ with effective radii below one parsec and enhanced surface densities compared to local analogues. These findings support the idea that extremely dense clusters, such as those used in our models, could plausibly form in the early Universe. However, our models do not include residual gas, which could remain trapped in such compact systems and influence early dynamical evolution and VMS formation.
@@
The presence of such dense clusters may also be required to explain chemical signatures observed in some high-redshift systems. Notably, the candidate galaxy GN-z11 at $z=10.6$ (Bouwens et al., 2010; Tacchella et al., 2023; Nagele & Umeda, 2023; Maiolino et al., 2024) exhibits an unusually high nitrogen-to-oxygen abundance ratio, with $\log_{10}(\mathrm{N}/\mathrm{O})>-0.25$—a value roughly four times the solar ratio (Cameron et al., 2023). Similar enhancements have been reported in other high-$z$ sources (Marques-Chaves et al., 2023). Rapidly rotating Pop-III stars have been suggested as a potential source of nitrogen enrichment (Tsiatsiou et al., 2024; Nandal et al., 2024). Additionally, high nitrogen abundances could originate from hydrogen-burning nucleosynthesis in VMS formed via runaway stellar collisions (Charbonnel et al., 2023). Our results, which show rapid VMS growth within a few Myr, support this scenario by demonstrating that VMSs can form early and potentially enrich their surroundings before core collapse.
@@
Compact massive clusters may also form during major galaxy mergers, particularly at high redshift where galaxies are more gas-rich and metal-poor. Mergers can trigger large-scale inflows of gas, leading to starbursts and the formation of dense stellar systems containing $>10^{4}$ to $10^{6}$ stars (Renaud et al., 2015). While nearby merging systems like the Antennae galaxies have near-solar metallicities (Fall et al., 2005), similar processes occurring in early galaxies (e.g. Pettini et al., 1997) could create the ideal conditions for VMS formation. This has been demonstrated in simulations by Lahén et al. (2019, 2020b, 2020a), where massive clusters with masses $\gtrsim 10^{5}\leavevmode\nobreak\ \rm M_{\odot}$ formed in the aftermath of gas-rich dwarf galaxy mergers. These simulated clusters are broadly consistent with the initial conditions explored in our models, reinforcing the relevance of our results for understanding early IMBH seed formation in the cosmological context.
@@ 5.3 Outlook: Observational Signatures and Future Tests
Tidal disruption events (TDEs) represent a promising observational signature of massive black holes (BHs), including those that may have formed via the collisional pathways modeled in this work. These events occur when a star passes sufficiently close to a BH to be partially or fully disrupted and accreted, producing luminous flares across the electromagnetic spectrum (Hills, 1975; Rees, 1988). TDEs are associated with the so-called “loss cone”—a region of phase space containing stellar orbits that bring stars close enough to be disrupted (Chandrasekhar, 1942). In spherical systems, this region is rapidly depleted, and its replenishment through two-body relaxation is typically slow (Broggi et al., 2024). Although TDEs can contribute to BH growth, our simulations conclude shortly after IMBH formation, and longer-term dynamical evolution—including loss cone refilling—remains a key area for future study.
@@
In the “empty loss cone” regime, where stellar orbits rarely intersect the BH’s tidal radius, BH growth stalls unless aided by additional processes such as gas accretion or compact object mergers (e.g. Stone & Metzger, 2016; Ryu et al., 2020c, b, a; Stone et al., 2020). One particularly intriguing channel involves the gradual inspiral of compact remnants (neutron stars or stellar-mass BHs) around an IMBH—a class of events known as extreme mass ratio inspirals (EMRIs) (Hils & Bender, 1995). EMRIs are shaped by dynamical friction, relativistic precession, and tidal forces, and they represent a key gravitational wave (GW) signature of IMBHs in dense star clusters (Broggi et al., 2022). Our results, which demonstrate early IMBH formation via runaway collisions, suggest that such GW sources could appear much earlier in cosmic history than traditionally assumed.
@@
Crucially, these processes are not only theoretically predicted—they are increasingly accessible to observation. TDEs produce luminous flares observable with facilities such as the Hubble Space Telescope (HST) (Leloudas et al., 2016) and the Neil Gehrels Swift Observatory (Swift) (Brown et al., 2016). Meanwhile, mergers involving compact objects—including EMRIs—are detectable with current GW observatories such as LIGO, Virgo, and KAGRA (Abbott et al., 2024), and will be prime targets for future space-based missions like LISA (Amaro-Seoane et al., 2017; McCaffrey et al., 2025) and the Einstein Telescope (ET) (Punturo et al., 2010).
@@
Complementary electromagnetic observations will further aid in testing the predictions of our collisional IMBH formation scenario. The upcoming JASMINE mission will offer ultra-precise astrometry of the Milky Way’s nuclear star cluster, enabling dynamical BH mass measurements. The Extremely Large Telescope (ELT), with its near-infrared imager MICADO, will be capable of resolving the sphere of influence of BHs at distances up to twice that of JWST (Davies et al., 2018). Already, JWST has revealed a population of compact, red sources—the so-called ”Little Red Dots”—which are interpreted as dusty supermassive BHs (SMBHs) or obscured, star-forming galaxies (Kokorev et al., 2024; Akins et al., 2025; Matthee et al., 2024; Napolitano et al., 2024). These high-redshift objects exhibit unusually large BH-to-stellar mass ratios, up to two orders of magnitude higher than those seen locally (Goulding et al., 2023; Scoggins et al., 2023; Übler et al., 2024; Furtak et al., 2024; Juodžbalis et al., 2024). Such findings suggest the presence of ”overmassive” BHs in the early Universe and challenge conventional models of BH seed formation.
@@ Acknowledgements.
MCV acknowledges funding through ANID (Doctorado acuerdo bilateral DAAD/62210038) and DAAD (funding program number 57600326). Furthermore, acknowledges to the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. AWHK and RS acknowledge NAOC International Cooperation Office for its support in 2023 and 2024. RS acknowledges the support by the National Science Foundation of China (NSFC) under grant No. 12473017. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). NH is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). NH received financial support from the European Union’s HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Sklodowska-Curie grant agreement number 101086388 - Project acronym: LACEGAL. FFD and RS acknowledge support by the German Science Foundation (DFG, project Sp 345/24-1). PB thanks the support from the special program of the Polish Academy of Sciences and the U.S. National Academy of Sciences under the Long-term program to support Ukrainian research teams grant No.  PAN.BFB.S.BWZ.329.022.2023. The work of PB was also supported by the grant No. BR24992759 “Development of the concept for the first Kazakhstan orbital cislunar telescope - Phase I”, financed by the Ministry of Science and Higher Education of the Republic of Kazakhstan. MAS acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101025436 (project GRACE-BH, PI: Manuel Arca Sedda). MAS acknowledge financial support from the MERAC foundation. AT is supported by Grants-in-Aid for Scientific Research (grant No. 19K03907 and 24K07040) from the Japan Society for the Promotion of Science. DRGS gratefully acknowledges support by the ANID BASAL project FB21003, as well as via Fondecyt Regular (project code 1201280) and ANID QUIMAL220002. DRGS thanks for funding via the Alexander von Humboldt - Foundation, Bonn, Germany. AH and MG were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191. Xiaoying Pang acknowledges the financial support of the National Natural Science Foundation of China through grants 12173029 and 12233013 TN acknowledges the support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 of the DFG Cluster of Excellence ”ORIGINS”. RC is supported in part by the National Key Research and Development Program of China and the Zhejiang provincial top level research support program. AE acknowledge financial support from the Center for Astrophysics and Associated Technologies CATA (FB210003). Computations were performed on the HPC system Raven at the Max Planck Computing and Data Facility, and we also acknowledge the Gauss Centre for Supercomputing e.V. for computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC).
Hardware settings
The simulations are executed on the Raven high performance computing system at the Max Planck Computing and Data Facility (MPCDF) [^3] Raven: https://www.mpcdf.mpg.de/services/supercomputing/raven. Here, we use from one to three full GPU computing node for the simulation, which gives the optimal performance and most cost-effective use of the computing resources. Each GPU computing node is equipped 4 Nvidia A100 GPUs (4 x 40 GB HBM2 memory per node and NVLink). Furthermore, each node has one CPU hosts, which is an Intel Xeon IceLake Platinum 8360Y with 72 CPU cores. The RAM is 512 GB per node or 256 GB of RAM depending on the node you get assigned to via the slurm system. For a detailed account including benchmarking concerning the optimal performance of the Nbody6++GPU code on hybrid, GPU-accelerated supercomputers, please refer to the recent review by Spurzem & Kamlah (2023) and sources therein. Even using this state-of-the-art high performance computing (HPC) infrastructure, the simulation take up to 2 months of real time to reach a star cluster age of 5 Myr. Tests and developments on machines such as juwels-booster in Germany  [^4] Juwels-booster: https://apps.fz-juelich.de/jsc/hps/juwels/booster-overview.html, Lumi in Finland and Leonardo in Italy were done. We expect more data from these infrastructures soon.
@@
The underlying data, including the initial model used in this work, as well as the output and diagnostic files from both the Nbody6++GPU and MOCCA simulations, are publicly available at https://doi.org/10.5281/zenodo.15283075. The Nbody6++GPU code version that includes the level C stellar evolution prescriptions (Kamlah et al., 2022a) is publicly available. The McLuster version used to generate the initial conditions is also publicly available as described in Leveque et al. (2022b, a).
@@ Appendix A Stellar evolution treatment
In our work, we follow the level C stellar evolution package, as presented in (see Table X in Kamlah et al. 2022a) and the stellar evolution fitting formulae from SSE by Hurley et al. (2000a), which also describes the stellar evolution routines and parameters in detail, with updates as described in Spurzem & Kamlah (2023) and in Sections below.
[…]
@@
For the compact objects evolution, we use remnant mass prescriptions following Fryer et al. (2012), and here we choose the delayed SNe mechanism as the slow extreme of the convection-enhanced, neutrino-driven, SNe paradigm. We use standard momentum conserving fallback-scaled kicks (drawn from a Maxwellian distribution with a dispersion of 265.0 $\mathrm{km\,s}^{-1}$ from Hobbs et al. (2005)) for the NSs and BHs (Belczynski et al. 2008). This is not the case for the NSs and BHs produced by the electron-capture SNe (ECSNe), accretion-induced collapse (AIC) and merger-induced collapse (MIC) (Podsiadlowski et al. 2004; Ivanova et al. 2008; Gessner & Janka 2018; Leung et al. 2020) and that are subject to low velocity kicks (drawn from a Maxwellian distribution with a dispersion of 3.0 $\mathrm{km\,s}^{-1}$ from Gessner & Janka (2018)).
@@
For the natal spins of the compact objects, we have (i) BHs receive natal spins following the Geneva models (Banerjee et al. 2020; Banerjee 2021). and (ii) WDs receive natal kicks following Fellhauer et al. (2003) (drawn from a Maxwellian distribution with a dispersion of 2.0 $\mathrm{km\,s}^{-1}$ but, which is capped at 6.0 $\mathrm{km\,s}^{-1}$). We switch on the weak (pulsational) pair instability SNe ((P)PISNe) following Leung & Nomoto (2019).
@@
Finally (iii) the NS did not receive any upgrade, following (Hurley et al. 2000a), where the spin is derived from the classical polytrope equation for n=3/2. Lastly, we briefly outline the settings concerning the runs with general relativity merger recoil kicks. When there is a collision between two compact objects (in these simulations there will be NSs or BHs), then the merger product will have a spin drawn from a Maxwellian distribution with $\sigma=0.2$. In the collision itself, there is a mass loss due to the emission of GWs. Therefore, the final merger product will have 0.985 times the sum of the masses of the compact objects that participated in the merger. The kick velocity is then calculated as described in Arca Sedda et al. (2023a). We stress that these kicks apply to all compact object mergers, so also compact object binary mergers consisting of WDs and NSs.

### v2

@@ 5.2 Observational context for VMS formation at high redshift
Observing massive, compact, and low-metallicity star clusters around the time of core collapse remains extremely challenging. However, high-redshift galaxies are widely expected to form such clusters efficiently due to their high gas fractions and low mlicities (e.g. Tacconi2010), as well as their clumpy substructures that favor gravitational fragmentation and rapid star formation (e.g. Devereaux2024). These conditions closely mirror those adopted in our simulations, reinforcing the plausibility of early VMS formation via runaway stellar collisions. We note that the majority of massive stars observed have masses of approximately $100-300\penalty 10000\ \rm M_{\odot}$ (Doran2013; Keszthelyi2025), stars with masses above this limit remain purely theoretical.
Recent observations using strong gravitational lensing have begun to probe star-forming clumps and stellar clusters in high-redshift galaxies. For example, Claeyssens2023 analyzed clump populations in galaxies at z=1z=18.58.5, identifying half-mass radii ranging from a few to several tens of parsecs. They also found evidence for a mass–size scaling relation that may lie above the local relation (Brown2021), although the observed population is likely biased toward more massive and extended systems. Nevertheless, extremely compact clusters are still expected to form in high-density environments, even if they remain observationally elusive.
@@
From a theoretical standpoint, the Toomre instability criterion implies that the typical cluster size should scale inversely with the square root of the gas surface density, i.e., $R_{h}\propto\Sigma_{\mathrm{gas}}^{-1/2}$ (Toomre1964). This suggests that star clusters formed in gas-rich, high-redshift galaxies should be significantly more compact than their local counterparts—a trend also supported by simulations and nearby observations (Choksi2021). Reinforcing this, Adamo2024 recently reported several compact clusters at $z\sim 10$ with effective radii below one parsec and enhanced surface densities compared to local analogues. These findings support the idea that extremely dense clusters, such as those used in our models, could plausibly form in the early Universe. However, our models do not include residual gas, which could remain trapped in such compact systems and influence early dynamical evolution and VMS formation.
@@
The presence of such dense clusters may also be required to explain chemical signatures observed in some high-redshift systems. Notably, the candidate galaxy GN-z11 at $z=10.6$ (Bouwens2010; Tacchella2023; Nagele2023; Maiolino2024) exhibits an unusually high nitrogen-to-oxygen abundance ratio, with $\log_{10}(\mathrm{N}/\mathrm{O})>-0.25$—a value roughly four times the solar ratio (Cameron2023). Similar enhancements have been reported in other high-$z$ sources (Marques-Chaves2024). Rapidly rotating Pop-III stars have been suggested as a potential source of nitrogen enrichment (Tsiatsiou2024; Nandal2024). Additionally, high nitrogen abundances could originate from hydrogen-burning nucleosynthesis in VMS formed via runaway stellar collisions (Charbonnel2023). Our results, which show VMS growth within a few Myr support this scenario by demonstrating that VMSs can form early and potentially enrich their surroundings before core collapse.
@@
Compact massive clusters may also form during major galaxy mergers, particularly at high redshift where galaxies are more gas-rich and m-poor. Mergers can trigger large-scale inflows of gas, leading to starbursts and the formation of dense stellar systems containing $10^{4}$ to $10^{6}$ stars (Renaud2015). While nearby merging systems like the Antennae galaxies have near-solar mlicities (Fall2005), similar processes occurring in early galaxies (e.g. Pettini1997) could create the ideal conditions for VMS formation. This has been demonstrated in simulations by Lahe2019; Lahe2020a; Lahe2020, where massive clusters with masses $\gtrsim 10^{5}\penalty 10000\ \rm M_{\odot}$ formed in the aftermath of gas-rich dwarf galaxy mergers. These simulated clusters are broadly consistent with the initial conditions explored in our models, reinforcing the relevance of our results for understanding early IMBH seed formation in the cosmological context.
@@ 5.3 Outlook: Observational Signatures and Future Tests
Tidal disruption events (TDEs) represent a promising observational signature of massive black holes (BHs), including those that may have formed via the collisional pathways modeled in this work. These events occur when a star passes sufficiently close to a BH to be partially or fully disrupted and accreted, producing luminous flares across the electromagnetic spectrum (Hils1975; Rees1988). TDEs are associated with the so-called “loss cone”—a region of phase space containing stellar orbits that bring stars close enough to be disrupted (Chandrasekhar1942). In spherical systems, this region is rapidly depleted, and its replenishment through two-body relaxation is typically slow (Broggi2024). Although TDEs can contribute to BH growth, our simulations conclude shortly after IMBH formation, and longer-term dynamical evolution—including loss cone refilling—remains a key area for future study.
@@
In the “empty loss cone” regime, where stellar orbits rarely intersect the BH’s tidal radius, BH growth stalls unless aided by additional processes such as gas accretion or compact object mergers (e.g. Stone2016; Ryu2020a; Ryu2020b; Ryu2020; Stone2020). One particularly intriguing channel involves the gradual inspiral of compact remnants (neutron stars or stellar-mass BHs) around an IMBH—a class of events known as extreme mass ratio inspirals (EMRIs) (Hils1995). EMRIs are shaped by dynamical friction, relativistic precession, and tidal forces, and they represent a key gravitational wave (GW) signature of IMBHs in dense star clusters (Broggi2022). Our results, which demonstrate early IMBH formation via runaway collisions, suggest that such GW sources could appear much earlier in cosmic history than traditionally assumed.
@@
Crucially, these processes are not only theoretically predicted—they are increasingly accessible to observation. TDEs produce luminous flares observable with facilities such as the Hubble Space Telescope (HST) (Leloudas2016) and the Neil Gehrels Swift Observatory (Swift) (Brown2016). Meanwhile, mergers involving compact objects—including EMRIs—are detectable with current GW observatories such as LIGO, Virgo, and KAGRA (Abbott2024), and will be prime targets for future space-based missions like LISA (Amaro-Seoane2017; McCaffrey2025) and the Einstein Telescope (ET) (Punturo2010).
@@
Complementary electromagnetic observations will further aid in testing the predictions of our collisional IMBH formation scenario. The upcoming JASMINE mission will offer ultra-precise astrometry of the Milky Way’s nuclear star cluster, enabling dynamical BH mass measurements. The Extremely Large Telescope (ELT), with its near-infrared imager MICADO, will be capable of resolving the sphere of influence of BHs at distances up to twice that of JWST (Davies2018). Already, JWST has revealed a population of compact, red sources—the so-called ”Little Red Dots”—which are interpreted as dusty supermassive BHs (SMBHs) or obscured, star-forming galaxies (Kokorev2024; Akins2025; Matthee2024; Napolitano2025). These high-redshift objects exhibit unusually large BH-to-stellar mass ratios, up to two orders of magnitude higher than those seen locally (Goulding2023; Scoggins2023; Ubler2024; Furtak2024; Juodzbalis2024). Such findings suggest the presence of ”overmassive” BHs in the early Universe and challenge conventional models of BH seed formation.
@@ Acknowledgements.
MCV acknowledges funding through ANID (Doctorado acuerdo bilateral DAAD/62210038) and DAAD (funding program number 57600326). Furthermore, acknowledges to the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. AWHK and RS acknowledge NAOC International Cooperation Office for its support in 2023, 2024, and 2025. RS acknowledges the support by the National Natural Science Foundation of China (NSFC) under grant No. 12473017. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). NH is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). NH received financial support from the European Union’s HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Sklodowska-Curie grant agreement number 101086388 - Project acronym: LACEGAL. This material is based upon work supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS. FFD and RS acknowledge support by the German Science Foundation (DFG, project Sp 345/24-1). PB thanks the support from the special program of the Polish Academy of Sciences and the U.S. National Academy of Sciences under the Long-term program to support Ukrainian research teams grant No. PAN.BFB.S.BWZ.329.022.2023. The work of PB was also supported by the grant No. BR24992759 “Development of the concept for the first Kazakhstan orbital cislunar telescope - Phase I”, financed by the Ministry of Science and Higher Education of the Republic of Kazakhstan. MAS acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101025436 (project GRACE-BH, PI: Manuel Arca Sedda). MAS acknowledge financial support from the MERAC foundation. AT is supported by Grants-in-Aid for Scientific Research (grant No. 19K03907 and 24K07040) from the Japan Society for the Promotion of Science. DRGS gratefully acknowledges support by the ANID BASAL project FB21003, as well as via Fondecyt Regular (project code 1201280) and ANID QUIMAL220002. DRGS thanks for funding via the Alexander von Humboldt - Foundation, Bonn, Germany. AH and MG were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191. Xiaoying Pang acknowledges the financial support of the National Natural Science Foundation of China through grants 12173029 and 12233013. TN acknowledges the support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 of the DFG Cluster of Excellence ”ORIGINS”. RC is supported in part by the National Key Research and Development Program of China and the Zhejiang provincial top level research support program. AE acknowledge financial support from the Center for Astrophysics and Associated Technologies CATA (FB210003). Computations were performed on the HPC system Raven at the Max Planck Computing and Data Facility, and we also acknowledge the Gauss Centre for Supercomputing e.V. for computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC). Finally, we also acknowledge A. Sander and his team for helpful comments.
@@
The underlying data, including the initial model used in this work, as well as the output and diagnostic files from both the Nbody6++GPU and MOCCA simulations, are publicly available at https://doi.org/10.5281/zenodo.15283075. The Nbody6++GPU code version that includes the level C stellar evolution prescriptions (Kamlah2022a) is publicly available. The McLuster version used to generate the initial conditions is also publicly available as described in Leveque2022a; leveque2022.
@@ Appendix A Stellar evolution treatment
In our work, we follow the level C stellar evolution package, as presented in (see Table X in Kamlah2022a) and the stellar evolution fitting formulae from SSE by Hurley2000, which also describes the stellar evolution routines and parameters in detail, with updates as described in Spurzem2023 and in Sections below.
We use the metallicity dependent winds following Vink2001; Vink2002; Vink2005; Belczynski2010 across the full mass range.
Here, we use the recipe for the stellar winds that does not take into account the bi-stability jump (Belczynski2010), and explained in the following. We highlight here that the winds of the massive (MS) stars have enormous effect on our results. The mass growth of the massive MS star via collisions is in constant competition with the mass loss driven by the strong, hot, fast OB-type winds for massive MS stars (Vink2021). We have a strong mass loss recipe for massive OB-type MS stars which takes into account the so-called bi-stability. The bi-stability jump causes a sudden jump in the line-driving characteristic of Iron. Specifically, this is the result of a recombination of the Fe(IV) to the Fe(III) ion at effective stellar temperatures of around $T\sim 25\,000$ K. The Fe(III) ion is a much more effective driver for radiative transport and therefore, the mass loss is enhanced (for more details see e.g. Belczynski2010; Vink2021; Bjorklund2023). This mass loss recipe can be chosen in the future, due to accumulating evidence that the bi-stability jump actually exists in nature (see e.g. recent observations by Bernini-Peron2023, and sources therein), even though the exact impact is still unexplored in star cluster dynamics.
@@
For the compact objects evolution, we use remnant mass prescriptions following Fryer2012, and here we choose the delayed SNe mechanism as the slow extreme of the convection-enhanced, neutrino-driven, SNe paradigm. We use standard momentum conserving fallback-scaled kicks (drawn from a Maxwellian distribution with a dispersion of 265.0 $\mathrm{km\,s}^{-1}$ from Hobbs2005) for the NSs and BHs (Belczynski2008). This is not the case for the NSs and BHs produced by the electron-capture SNe (ECSNe), accretion-induced collapse (AIC) and merger-induced collapse (MIC) (Podsiadlowski2004; Ivanova2008; Gessner2018; Leung2020) and that are subject to low velocity kicks (drawn from a Maxwellian distribution with a dispersion of 3.0 $\mathrm{km\,s}^{-1}$ from Gessner2018).
@@
For the natal spins of the compact objects, we have(i) BHs receive natal spins following the Geneva models (Banerjee2020; Banerjee2021) and (ii) WDs receive natal kicks following Fellhauer2003 (drawn from a Maxwellian distribution with a dispersion of 2.0 $\mathrm{km\,s}^{-1}$ but, which is capped at 6.0 $\mathrm{km\,s}^{-1}$). We switch on the weak (pulsational) pair instability SNe ((P)PISNe) following Leung2019.
@@
Finally (iii) the NS did not receive any upgrade, following (Hurley2000), where the spin is derived from the classical polytrope equation for n=3/2. Lastly, we briefly outline the settings concerning the runs with general relativity merger recoil kicks. When there is a collision between two compact objects (in these simulations there will be NSs or BHs), then the merger product will have a spin drawn from a Maxwellian distribution with $\sigma=0.2$. In the collision itself, there is a mass loss due to the emission of GWs. Therefore, the final merger product will have 0.985 times the sum of the masses of the compact objects that participated in the merger. The kick velocity is then calculated as described in ArcaSedda2023. We stress that these kicks apply to all compact object mergers, so also compact object binary mergers consisting of WDs and NSs.